OpenStudy (anonymous):

Find the tangent to the curve at the point Use f (x) = 2/x at x = 4

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

To find the tangent to the curve at the given point

OpenStudy (anonymous):

using this equation given the point at x=4 \[f \prime (a)=\lim_{x \rightarrow a} \frac{ f(x)-f(a) }{ x-a }\] substitute \[f \prime(4)=\lim_{x \rightarrow 4} \frac{ \frac{ 2 }{ x } -2/4}{ x- 4}\] then solve

OpenStudy (anonymous):

you get the slope of the tangent

OpenStudy (anonymous):

how would I solve for the slope of the tangent?

OpenStudy (anonymous):

get rid of the denominators by mutiplying the top and bottom by (x)(4)